# AfriDoc-MT: Document-level MT Corpus for African Languages

### v1

@@ Acknowledgements
This work was carried out with support from Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. This research project has benefitted from the Microsoft Accelerate Foundation Models Research (AFMR) grant program. We acknowledges the support of OpenAI for providing API credits through their Researcher Access API programme. Jesujoba Alabi acknowledges the support of the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Miaoran Zhang received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102. Lastly, we also appreciate the translators whose names we have listed below:
@@ Appendix A More details about AfriDoc-MT
Table 9 shows the average number of white-space-separated tokens for sentences across various domains and their corresponding translations in all the languages including English. The health domain has more tokens on average than tech. Hausa and Yorùbá have more tokens on average than English, possibly because they are descriptive languages, while Swahili has a comparably similar length to English. However, Amharic and Zulu have relatively shorter average lengths, demonstrating interesting linguistic properties.
A.1 Translation Guideline
Below is the translation guideline aside the details shared at the workshop on translation and terminology creation.
@@ A.1 Translation Guideline
| Domain | eng | amh | hau | swa | yor | zul |
| Sentence | ||||||
| health | ||||||
| tech | ||||||
| Document | ||||||
| health | ||||||
| tech | ||||||
A.2 Quality evaluation of the translations
As part of the human translation process, we conducted quality estimation to assess the translations. For this purpose, we used AfriCOMET101010https://huggingface.co/masakhane/africomet-qe-stl. Given a translated sentence in any African language and its corresponding source English sentence, AfriCOMET generates a score between 0 and 1, where 0 indicates poor quality and higher values signify better quality. The translators, in collaboration with the language coordinators, were tasked with reviewing instances that had quality estimation scores below 0.65. This step was essential to identify and correct low-quality translations.
Figure 5 illustrate the distribution of the final quality scores for the five languages and both domains. Our manuel check indicates that QE scores below 0.65 are not necessarily indicative of poor translations, which is consistent with the findings of Adelani et al. (2024b). We attribute this observation to factors such as domain shift, translation length, and other potential influences, which warrant further investigation in future research.
A.3 Creation of Pseudo-documents for AfriDoc-MT
Given that the translated documents vary in length in terms of sentences and tokens, and considering the maximum token length limitations of the different LLMs used, we adopted a chunking approach for document-level evaluation. In this approach, documents were divided into smaller pseudo-documents that fit within the maximum length constraints of the models. To establish an appropriate chunk size, each document was divided into fixed-size chunks of k sentences, with the possibility that the final chunk may contain fewer than k sentences. These sentence groups, referred to as pseudo-documents, were used for document-level translation.
We conducted an initial analysis, testing different values for k (5, 10, and 25), with k=1 serving as our sentence-level setup. Table 10 presents the resulting number of parallel pseudo-documents, as well as the average number of tokens per pseudo-document per language for the various model tokenizers, including the 95th percentile token count. Our analysis revealed that Amharic and Yorùbá —languages with unique characteristics such as non-Latin scripts and diacritics, respectively—had the largest average token counts across the tokenizers. Additionally, domain with highest number of average tokens for pseudo-document vary from language to language.
To accommodate both languages in our experiments, we chose pseudo-documents with k=10. However, for the SFT models described in Section Section B.2, we used both k=5 and k=10.
@@ B.1.1 Translation Models
M2M-100 Fan et al. (2020) is a transformer-based multilingual neural translation model from Meta, trained to translate between 100 languages, including several African languages. It has three variants of different sizes: 400M parameters, 1.2B parameters, and 12B parameters. For our experiment, we evaluated the 400M and 1.2B variants.
@@
Team et al. (2024) is a model similar to M2M-101, with broader coverage, trained to translate between just over 200 languages, including more than 50 African languages. It also has different sizes: 600M, 1.3B, 3.3B, and 54B parameters. However, for this work, we evaluated the first three variants.
@@
Kudugunta et al. (2023) is a multilingual translation model based on the T5 architecture, covering 450 languages, including many African languages. It was trained on data collected from the CommonCrawl dataset. The dataset underwent a thorough self-audit to filter out noisy content and ensure its quality for training machine translation models.
@@
Elmadany et al. (2024); Adebara et al. (2024) is another multilingual but african-centric translation model based on the T5 architecture, covering 150 language pairs of African languages. It was first pre-trained on large multilingual texts covering over 500 African languages and then finetuned on translation task covering over 100 language pairs.
@@ B.1.2 Large Language Models
Gemma Team et al. (2024) is a decoder-only LLM trained on billions of tokens sourced from the web. The training data primarily consists of English-language text, but it also includes code and mathematical content. While Gemma2 has an English-centric focus, it also possesses multilingual capabilities. We evaluate the base Gemma2 model with 9B parameters, as well as its instruction-tuned version.
@@
Dubey et al. (2024) is another decoder-only LLM trained on trillions of tokens across multiple languages. It was fine-tuned using existing instruction datasets as well as synthetically generated instruction data to create its instruction-tuned version. One advantage LLama3.1 has over other models is its context window of 128K tokens, the largest among all models considered in this work, making it particularly suitable for document-based tasks such as document-level translation. We evaluate the base LLama3.1 model with 8B parameters, as well as its instruction-tuned version.
@@
Lu et al. (2024) is a multilingual LLM built on the LLama3 with 8B parameters as its base. It was trained on 102 languages, including several African languages, through continued pretraining. Using an English instruction dataset (Alpaca), it was further fine-tuned to create LLaMAX3-Alpaca. We evaluated both models and compared their performance across various tasks.
@@ B.2 Supervised Finetuning
We perform supervised fine-tuning to tailor LLMs for translation tasks. For training sentence-level MT systems, we use all parallel sentences from AfriDoc-MT to construct the training set, enabling the LLMs to translate across multiple directions and domains. Following Zhu et al. (2024b), we augment the parallel data with translation instructions, which are randomly sampled from a predefined set of 31 MT instructions for each training example.111111We use the same instruction set as described in Zhu et al. (2024b). For training document-level MT systems, we follow the same process, but train on longer segments formed by concatenating multiple sentences. In finetuning, we use a learning rate of and an effective batch size of 64. Models are trained for only one epoch, as further training does not yield improvements and may even lead to performance degradation.
Similarly, we fine-tuned the 1.3B version of NLLB-200 for sentence and pseudo-document (with 10 sentences) translation using the Fairseq Ott et al. (2019) codebase. We used all the training examples from 30 language directions across both domains. The model was fine-tuned for 50K steps using a learning rate of , token batch size of 2048 and a gradient accumulation of 2. The checkpoint with the lowest validation loss was selected as the best model for evaluation.
@@ B.3 Evaluation setup
All the models were evaluated using different tools. For example, both the NLLB-200 and M2M-100 models were evaluated with the Fairseq codebase, while Toucan and MADLAD-400 were evaluated using the Hugging Face (HF) codebase. All other LLMs, including LLama3.1 (both instruction-tuned and SFT models), Gemma, and Aya-101, were evaluated using EleutherAI LM Evaluation Harness (lm-eval) tool (Biderman et al., 2024). In all cases, greedy decoding was used.
All the models evaluated have different context lengths. For encoder-decoder models, M2M-100 and NLLB have a maximum sequence length of 1024 and 512 respectively. Aya-101 and MADALAD, based on the T5 architecture, do not have a pre-specified maximum sequence length, so we fixed their maximum sequence length to 1024 for all experiments involving encoder-decoder models. However, for decoder-only models, Gemma and LlamaX3 (based on LLama3) have a maximum sequence length of 8192, while LLama3.1 has a maximum sequence length of 128K. Since all the decoder-only models were evaluated using LM Eval, we used a similar setup for them, selecting the maximum length based on the specific needs of each model.
@@
Table 11 shows the maximum number of generation tokens we set when translating between English and African languages. These numbers were chosen based on the statistics from Table 10. However, for Amharic, when translating pseudo-documents with 25 sentences and full documents, there were instances exceeding the 95th percentile derived from the training statistics. Therefore, we increased the token limit specifically for Amharic.
@@ B.4 Evaluation prompts
While the translation models we evaluated require no prompts, MADLAD-400, on the other hand, requires a prefix of the form <2xx> token, which is prepended to the source sentence. Here, xx indicates the target language using its language code (e.g., "sw" for Swahili). Similarly, Toucan uses just the target language ISO-693 code as prefix, which is prepended to the source sentence (e.g., "swa" for Swahili). For other models, including Aya-101, we used three different prompts for sentence-level translation and document translation experiments. The main difference between the prompts for these tasks is the explicit mention of "text" or "document" within the prompt, as shown in Table 19. For the base models Gemma2, Llama3.1, LLaMAX3, and Aya-101, we prompted them directly using the respective prompts. However, for the instruction-tuned versions of Gemma2 and Llama3.1, we used their respective chat templates. For all Alpaca-based models, including our SFT models, we used the Alpaca template.
| Model | Setup | eng X | X eng | ||||||||
| d-CHRF | Fluency | CE | LE | GE | d-CHRF | Fluency | CE | LE | GE | ||
| Aya-101 | Sent | ||||||||||
| Doc10 | |||||||||||
| GPT-3.5 | Sent | ||||||||||
| Doc10 | |||||||||||
| LLaMAX3-SFT1 | Sent | ||||||||||
| Doc10 | |||||||||||
| LLaMAX3-SFT10 | Doc10 | ||||||||||
| Sent | |||||||||||
| GPT-4o | Doc10 | ||||||||||
@@ B.5 Evaluation metrics
We evaluate translation quality with BLEU Papineni et al. (2002) and CHRF Popović (2015) using SacreBLEU121212case:mixed|eff:no| tok:13a|smooth:exp|v:2.3.1, Post (2018). We run significance tests using bootstrap resampling and report the confidence interval for the scores, based on a sample size of . We also use AfriCOMET131313https://huggingface.co/masakhane/africomet-stl-1.1 (Wang et al., 2024a) to evaluate the quality of the translation outputs. We report the CHRF scores of the best prompt for each model and language direction in the main paper, with all additional results provided in the Appendix C. For document-level experiments, we evaluated the LLMs using the same three prompts as in the sentence-level experiment. For evaluation, we used BLEU and CHRF scores but excluded AfriCOMET due to its backbone model, AfroXLM-R-L Alabi et al. (2022); Adelani et al. (2024a), having a context length of 512 tokens. This made it impractical to compute COMET scores for document-level outputs.
@@ B.6 GPT-4o as an evaluator for machine translation
As a proxy for human evaluation, we use GPT-4o to assess the quality of translation output, as demonstrated by Sun et al. (2024), which shows a correlation with human judgment. Due to the cost of this task, we limited our evaluation to a few selected models, including Aya-101, GPT-3.5, GPT-4o, and LLaMAX3 fine-tuned on AfriDoc-MT sentences and pseudo-documents of 10 sentences. We compared translations performed at the sentence level and pseudo-document level in terms of fluency, content errors, and cohesion errors—specifically lexical (LE) and grammatical (GE) errors—using the same definitions as Sun et al. (2024).
@@
Fluency can only have values between 1 and 5; however, the other metrics, including CE, GE, and LE, do not have a specific range and can take on any value because they are counts. Refer to Sun et al. (2024) for more details about these metrics.
@@ C.1 Sentence-level evaluation
Given that AfriDoc-MT is a document-level translation dataset, and due to the limited context length of most translation models and LLMs, which makes it impossible to translate a full document at once, we opted to translate the sentences within the documents and then merge them back to form the complete document. This also serves as a baseline for document-level translation. In the main paper, we present the results for the best prompt for each language pair and model using d-CHRF. In this section, we also provide the full results on the merged documents using d-CHRF and d-BLEU in Tables 13 and 14. Furthermore, we present results for evaluating just the sentences (without merging them back into documents) using s-BLEU, s-CHRF, and s-COMET in Tables 15 and 16. Also, in Figures 17, 18, 19 and 20, we provide plots that summarizes some of the results in the table for few models. Although the main findings are summarized in the main draft, below are some other points we identify.
@@
Both versions of M2M-100, which was once a state-of-the-art translation model, are not competitive with other translation models such as Toucan, NLLB-200, and MADLAD-400, even when compared to models of similar sizes, across all metrics at both the sentence and document levels.
@@
Base LLMs without instruction tuning and supervised fine-tuning, such as Gemma2 and LLaMAX3, also do not show competitive translation performance. This can be explained by the fact that they are just language models with limited coverage of African languages. However, LLaMAX3, which was trained on more than 100 languages, including African languages, through continued pre-training, shows improved performance, surpassing LLama3.1-IT.

### v2

@@ Acknowledgements
This work was carried out with support from Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program. We acknowledge the support of OpenAI for providing API credits through their Researcher Access API programme. Jesujoba Alabi acknowledges the support of the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. The work has also been partially funded by BMBF through the TRAILS project (grant number 01IW24005). Miaoran Zhang received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102. CEB acknowledges her AI4S fellowship within the “Generación D” initiative by Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by NextGenerationEU through PRTR. R. Bawden’s participation was partly funded by her chair in the PRAIRIE institute, funded by the French national agency ANR, as part of the “Investissements d’avenir” programme under the reference ANR-19-P3IA-0001 and by the French Agence Nationale de la Recherche (ANR) under the projects TraLaLaM (“ANR-23-IAS1-0006”) and MaTOS (“ANR-22-CE23-0033”). David Adelani acknowledges the funding of IVADO and the Canada First Research Excellence Fund.
Also, we also appreciate the translators whose names we have listed below:
@@
Lastly, we thank the human annotators who were not part of the translation team.
@@ Appendix A More details about AfriDoc-MT
A.1 Translation guidelines
The translation guidelines, aside the details shared at the workshop on translation and terminology creation can be found below.
@@ A.1 Translation guidelines
A.2 Creation of pseudo-documents for AfriDoc-MT
Given that the translated documents vary in length in terms of sentences and tokens, and considering the maximum token length limitations of the different LLMs used, we adopted a chunking approach for document-level evaluation. In this approach, documents were divided into smaller pseudo-documents that fit within the maximum length constraints of the models. To establish an appropriate chunk size, each document was divided into fixed-size chunks of sentences, with the possibility that the final chunk may contain fewer than sentences. These sentence groups, referred to as pseudo-documents, were used for document-level translation.
We conducted an initial analysis, testing different values for (5, 10, and 25), with =1 serving as our sentence-level setup. Table˜11 presents the resulting number of parallel pseudo-documents, as well as the average number of tokens per pseudo-document per language for the various model tokenizers, including the 95th percentile token count. Our analysis revealed that Amharic and Yorùbá —languages with unique characteristics such as non-Latin scripts and diacritics, respectively—had the largest average token counts across the tokenizers. Additionally, the domain with the highest number of average tokens for pseudo-document varies from language to language.
To accommodate both languages in our experiments, we chose pseudo-documents with =10. However, for the SFT models described in Section˜B.2, we used both =5 and =10.
@@ B.1.1 Translation Models
M2M-100 (Fan et al., 2020) is a transformer-based multilingual neural translation model from Meta, trained to translate between 100 languages, including several African languages. It has three variants of different sizes: 400M parameters, 1.2B parameters, and 12B parameters. For our experiments, we evaluated the 400M and 1.2B variants.
@@
(NLLB Team et al., 2024) is a model similar to M2M-100, with broader coverage, trained to translate between just over 200 languages, including more than 50 African languages. It also has different sizes: 600M, 1.3B, 3.3B, and 54B parameters. For this work, we evaluated the first three variants.
@@
(Kudugunta et al., 2023) is a multilingual translation model based on the T5 architecture (Raffel et al., 2020), covering 450 languages, including many African languages. It was trained on data collected from the Common Crawl dataset. The dataset underwent a thorough self-audit to filter out noisy content and ensure its quality for training MT models.
@@
(Elmadany et al., 2024; Adebara et al., 2024) is another multilingual but Afro-centric translation model based on the T5 architecture, covering 150 language pairs of African languages. It was first pre-trained on large multilingual texts covering over 500 African languages and then finetuned on translation task covering over 100 language pairs.
@@ B.1.2 Large Language Models
(Gemma Team et al., 2024) is a decoder-only LLM trained on billions of tokens sourced from the web. The training data primarily consists of English-language text, but it also includes code and mathematical content. While Gemma2 has an English-centric focus, it also possesses multilingual capabilities. We evaluate the base Gemma2 model with 9B parameters, as well as its instruction-tuned version.
@@
(Dubey et al., 2024) is another decoder-only LLM trained on trillions of tokens across multiple languages. It was fine-tuned using existing instruction datasets as well as synthetically generated instruction data to create its instruction-tuned version. One advantage LLama3.1 has over other models is its context window of 128K tokens, the largest among all models considered in this work, making it particularly suitable for document-based tasks such as document-level translation. We evaluate the base LLama3.1 model with 8B parameters, as well as its instruction-tuned version.
@@
(Lu et al., 2024) is a multilingual LLM built on the LLama3 with 8B parameters as its base. It was trained on 102 languages, including several African languages, through continued pretraining. Using an English instruction dataset (Alpaca), it was further fine-tuned to create LLaMAX3-Alpaca. We evaluated both models and compared their performance across various tasks.
@@ B.2 Supervised Finetuning
We perform supervised fine-tuning to tailor LLMs for translation tasks. To train sentence-level MT systems, we use all parallel sentences from AfriDoc-MT to construct the training set, enabling the LLMs to translate across multiple directions and domains. Following Zhu et al. (2024b), we augment the parallel data with translation instructions, which are randomly sampled from a predefined set of 31 MT instructions for each training example.121212We use the same instruction set as described in (Zhu et al., 2024b). To train document-level MT systems, we follow the same process, but train on longer segments formed by concatenating multiple sentences. When fine-tuning, we use a learning rate of and an effective batch size of 64. Models are trained for only one epoch, as further training does not result in improvements and may even lead to performance degradation.
Similarly, we fine-tuned the 1.3B version of NLLB-200 for sentence and pseudo-document (with 10 sentences) translation using the Fairseq (Ott et al., 2019) codebase. We used all the training examples from 30 language directions across both domains. The model was fine-tuned for 50k steps using a learning rate of , token batch size of 2048 and a gradient accumulation of 2. The checkpoint with the lowest validation loss was selected as the best model for evaluation.
@@ B.3 Evaluation setup
The models were evaluated using different tools. For example, both the NLLB-200 and M2M-100 models were evaluated with the Fairseq codebase, while Toucan and MADLAD-400 were evaluated using the Hugging Face (HF) codebase. All other LLMs, including LLama3.1 (both instruction-tuned and SFT models), Gemma, and Aya-101, were evaluated using EleutherAI LM Evaluation Harness (lm-eval) tool (Biderman et al., 2024). In all cases, greedy decoding was used.
The models evaluated have different context lengths. For encoder-decoder models, M2M-100 and NLLB have a maximum sequence length of 1024 and 512 respectively. Aya-101 and MADALAD, based on the T5 architecture, do not have a pre-specified maximum sequence length, so we fixed their maximum sequence length to 1024 for all experiments involving encoder-decoder models. However, for decoder-only models, Gemma and LLaMAX3 (based on LLama3) have a maximum sequence length of 8192, while LLama3.1 has a maximum sequence length of 128K. Since all the decoder-only models were evaluated using LM Evaluation Harness, we used a similar setup for them, selecting the maximum length based on the specific needs of each model.
@@
Table˜12 shows the maximum number of generation tokens we set when translating between English and African languages. These numbers were chosen based on the statistics from Table˜11. However, for Amharic, when translating pseudo-documents with 25 sentences and full documents, there were instances exceeding the 95th percentile derived from the training statistics. Therefore, we increased the token limit specifically for Amharic.
@@ B.4 Evaluation prompts
While the translation models we evaluated do not require prompts, MADLAD-400, requires a prefix of the form <2xx> token, which is prepended to the source sentence. Here, xx indicates the target language using its language code (e.g., “sw” for Swahili). Similarly, Toucan uses just the target language ISO-693 code as prefix, which is prepended to the source sentence (e.g., “swa” for Swahili). For other models, including Aya-101, we used three different prompts for sentence-level translation and document translation experiments. The main difference between the prompts for these tasks is the explicit mention of “text” or “document” within the prompt, as shown in Table˜23. For the base models Gemma2, Llama3.1, LLaMAX3, and Aya-101, we prompted them directly using the respective prompts. However, for the instruction-tuned versions of Gemma2 and Llama3.1, we used their respective chat templates. For all Alpaca-based models, including our SFT models, we used the Alpaca template.
@@ B.5 Evaluation metrics
We evaluate translation quality with BLEU (Papineni et al., 2002) and ChrF (Popović, 2015) using SacreBLEU131313case:mixed|eff:no| tok:13a|smooth:exp|v:2.3.1, (Post, 2018). We run significance tests using bootstrap resampling and report the confidence interval for the scores, based on a sample size of . We also use AfriCOMET141414https://huggingface.co/masakhane/africomet-stl-1.1 (Wang et al., 2024a) to evaluate the quality of the translation outputs. We report the chrF scores of the best prompt for each model and language direction in the main paper, with all additional results provided in the Appendix˜C. For document-level experiments, we evaluated the LLMs using the same three prompts as in the sentence-level experiment. For evaluation, we used BLEU and chrF scores but excluded AfriCOMET due to its backbone model, AfroXLM-R-L (Alabi et al., 2022; Adelani et al., 2024a), having a context length of 512 tokens. This made it impractical to compute COMET scores for document-level outputs.
@@ B.6 GPT-4o as an evaluator for machine translation
We use GPT-4o to assess the quality of translation output, as demonstrated by Sun et al. (2025), which shows a correlation with human judgment. Due to the cost of this task, we limited our evaluation to a few selected models, including Aya-101, GPT-3.5, GPT-4o, and LLaMAX3 fine-tuned on AfriDoc-MT sentences and pseudo-documents of 10 sentences. We compared translations performed at the sentence-level and pseudo-document level in terms of fluency, content errors, and cohesion errors—specifically lexical (LE) and grammatical (GE) errors—using the same definitions as Sun et al. (2025).
@@
Fluency can only have values between 1 and 5. However, the other metrics, including CE, GE, and LE, do not have a specific range and can take on any value because they are counts. Refer to (Sun et al., 2025) for more details about these metrics.
B.7 Human Evaluation Setup
Beyond using GPT-4o as a judge, we also conduct human evaluation on a subset of outputs from GPT-3.5, LLaMAX3-SFT1, and LLaMAX3-SFT10 for two domains, focusing specifically on translation into five African languages due to cost constraints. Translation into English was excluded, as existing automatic metrics, including GPT-based evaluations, are already reliable for this direction.
For the human evaluation, three native speakers of the African languages—primarily translators involved in the dataset creation—were recruited. Each annotator was assigned 80 documents to evaluate, tasked with marking as many error spans as possible and rating the overall quality on a scale from 0 to 100. This annotation followed the error span annotation (ESA) (Kocmi et al., 2024) protocol as implemented within the Appraise Evaluation Framework (Federmann, 2018). To assess consistency and inter-annotator agreement, 30 of the 80 documents were shared among all three annotators. Table˜13 shows statistics for 80 documents sampled from the models in both domains for each annotator. Each annotator was remunerated with $55.151515Annotation protocol.
| Model | Setup | Full | Shared | ||
| health | tech | health | tech | ||
| GPT-3.5 | Sent. | 5 | 5 | - | 5 |
| Pseudo. | 5 | 5 | - | 5 | |
| LLaMAX3-SFT1 | Sent. | 5 | 5 | 5 | - |
| Pseudo. | 5 | 5 | 5 | - | |
| LLaMAX3-SFT10 | Pseudo | 5 | 5 | 5 | 5 |
| Total | 25 | 25 | 15 | 15 | |
B.8 Qualitative Analysis
Alongside the human direct assessment of the translation outputs, we shared a subset of the outputs with one author per language, each a native speaker. They were tasked with analyzing the outputs to answer two key questions: (1) What common errors or flaws do the models exhibit across different setups? and (2) How fluent are the translation outputs produced by the models across the various settings?
@@ C.1 Sentence-level evaluation
Given that AfriDoc-MT is a document-level translation dataset, and due to the limited context length of most translation models and LLMs, which makes it impossible to translate a full document at once, we opted to translate the sentences within the documents and then merge them back to form the complete document. This also serves as a baseline for document-level translation. In the main paper, we present the results for the best prompt for each language pair and model using d-chrF. In this section, we also provide the full results on the merged documents using d-chrF and d-BLEU in Tables˜15 and 16. Furthermore, we present results for evaluating just the sentences (without merging them back into documents) using s-BLEU, s-chrF, and s-COMET in Tables˜17 and 18. In Figures˜16, 17, 18 and 19, we provide plots that summarize some of the results in the table for a few models. Although the main findings are summarized in the main draft, below are some other points we identify.
@@
Neither version of M2M-100, which was once a state-of-the-art translation model, is competitive with other translation models such as Toucan, NLLB-200, and MADLAD-400, even when compared to models of similar sizes, across all metrics at both the sentence and document levels.
@@
Base LLMs without instruction tuning and supervised fine-tuning, such as Gemma2 and LLaMAX3, do not show competitive translation performance either. This can be explained by the fact that they are just language models with limited coverage of African languages. However, LLaMAX3, which was trained on more than 100 languages, including African languages, through continued pre-training, shows improved performance, surpassing LLama3.1-IT.